OpenStudy (austinl):

Solve the given initial value problem and determine at least approximately where the solution is valid. \((9x^2+y-1)-(4y-x)y\prime=0,~~~y(1)=0\)

OpenStudy (austinl):

I already know that this differential equation is not exact, so, how to solve from there.

OpenStudy (psymon):

Did the question say it's not exact? I can rewrite it by factoring out a negative in the (4y-x) portion. \[(9x ^{2}+y-1)dx+(x-4y)dy = 0\] \[\frac{ \delta M }{ \delta y }(9x ^{2}+y-1) = 1\] \[\frac{ \delta N }{ \delta x }(x-4y) = 1\]

OpenStudy (austinl):

I know where I went wrong, I had: \(\dfrac{\partial N}{dx}(4y-x)\) I forgot to distribute the negative to get, \(\dfrac{\partial N}{dx}(x-4y)\)

OpenStudy (psymon):

Bingo. Then just solve like normal, it works out perfect.

OpenStudy (austinl):

Very cool, thanks for letting me know my silly error :P

OpenStudy (psymon):

No problem :3

OpenStudy (austinl):

Would I use the initial value portion of the problem to solve for C or h(y)?

OpenStudy (psymon):

You dont need the initial value portion to get h(y).

OpenStudy (austinl):

Oh, derp. I feel really really silly right now.

OpenStudy (psymon):

Mhm. Youre just doing a bunch of flip-flopping with which variables you're giving respect to. You do M with respect to y to check for exactness then integrate with respect to x to get h(y), then differentiatr with respect to y again and set that result = to Ny. So we would have: \[\int\limits_{}^{}(9x ^{2}+y-1)dx = 3x ^{3}+xy - x + h(y) \]Differentiate with respect to y: \[\frac{ \delta M }{ \delta y }(3x ^{3}+xy-x+h(y)) = x + h'(y) \]Now this value gets set to N, so we have: \[x+h'(y) = -4y + x \implies h'(y) = -4y\] Then integrate both sides to get h(y) \[\int\limits_{}^{}h'(y) = \int\limits_{}^{}-4ydy\] \[h(y) = -2y ^{2} \] Replace this result in the previous integration and set it all equal to C \[3x ^{3}+xy - x -2y ^{2}= C\] Now use the initial value to get C> If all of that makes sense, lol.

OpenStudy (austinl):

I have: \(C=3x^3-2y^2+x(y-1)=3x^2-2y^2+xy-1\) And now that I have this typed, I see you have the same answer.

OpenStudy (psymon):

Which is awesome, good job xD

OpenStudy (austinl):

Thank you, time for another silly question. \(y(1)=0\) What form is this? \(y(x)=?\)

OpenStudy (psymon):

Yeah, pretty much. When x is 1, y is 0 would be the interpretation.

OpenStudy (austinl):

\(2=3x^3+xy-x-2y^2\) Is this correct?

OpenStudy (psymon):

Thats what I got.

OpenStudy (austinl):

Wait, I need to solve for y....... oh my god.

OpenStudy (psymon):

I suppose you could, yeah, lol. Depends if you have to or not.

OpenStudy (austinl):

According to the back of my book I do..... because I also need to determine at least approximately where the solution is valid.

OpenStudy (psymon):

You know how to go about that then?

OpenStudy (austinl):

Now that I think about it no, my professor never discussed it and looking at the examples, none of them cover it.

OpenStudy (austinl):

Oh, and, \(y=\dfrac{x-\sqrt{24x^3+x^2-8x-16}}{4}\)

OpenStudy (psymon):

Lol, I didnt even have to write it out it seems xD Or did you look up the answer?

OpenStudy (austinl):

I typed it into my calculator and verified with my book. I wasn't even going to try with that one.

OpenStudy (psymon):

Well, the way you do it is you treat the x's like constants and then you can say that this is a quadratic in y, which allows you to use the quadratic formula: So just set everythign equal to 0 and treat it like y^2 + y + c using x's and numbers as constants. I have to go, though, sorry D: Try it out and see if you can get it. Oh, and at the end, the reason its x - instead of +/- is because you need to plug in the intial condition and see which sign makes it true, + or -. But yeah, good luck and glad I could help ^_^

OpenStudy (austinl):

How can you determine approximately where the solution is valid?

OpenStudy (austinl):

Oh well, thank you very very much!